OpenStudy (anonymous):

Choose the correct simplification of (6x − 5)(2x2 − 3x − 6) Ive been using the foil methd but cant get it?

OpenStudy (anonymous):

You can't use FOIL here, that's for quadratics, but what you can do is multiple 6x to all three of the second polynomial's terms and -5 to all three of the second polynomials terms. Make sense?

OpenStudy (anonymous):

You'll then combine like terms, so that means stuff like: 3x + 4x - 2x = 5x

OpenStudy (anonymous):

multiply* not multiple lol

OpenStudy (anonymous):

PS: Welcome to OpenStudy :-)

OpenStudy (anonymous):

so its like im distrubuting ?

OpenStudy (anonymous):

In a sense, yes

OpenStudy (anonymous):

i.e.: 6x * 2x\(^2\) = ?

OpenStudy (anonymous):

x^2 = x\(^2\) if you're not familiar with how to write exponents on computers :-)

OpenStudy (anonymous):

x\(^1\) = x so (6 * x) * (2 * x * x) = 6x * 2x\(^2\) = 12x\(^{1+2}\) = 12x\(^3\) Next, 6x * -3x = -18x\(^2\)

OpenStudy (anonymous):

Aright thankyou !(:

OpenStudy (anonymous):

that -18x^2 adds with -10x^2 (from -5 * 2x^2) to make -28x^2 here for instance: \(12 x^3-28 x^2-21 x+30\)

OpenStudy (anonymous):

Thankyou , in the course im taking they ony realy covered foil

OpenStudy (anonymous):

FOIL does the same thing actually, you just only have 2 pairs of terms per polynomial, so it's just a simple mnemonic just to help you memorize so you don't forget a term to multiply, nothing more, nothing less. :-)

OpenStudy (anonymous):

In this case you have a 2-term polynomial (aka. a binomial) and a 3-term polynomial (aka. a trinomial, or in this case specifically a quadratic)